OpenStudy (anonymous):

Which choice is the equation of a line that passes through point (7,3) and is parallel to the line represented by this equation? Y=2/7x-3 A)7x+3y=2 B)y=2/7x+1 C)y=-7/2x-3 D)2x+7y=1

OpenStudy (anonymous):

If the line is parallel, then the coefficients of \(x\) and \(y\) will be the same, so we have: \[ y=\frac 27x - b \]We plug in \((x,y) = (7,3)\) and get: \[ (3) = \frac 27(7) - b \]This equation allows us to solve for \(b\).

OpenStudy (anonymous):

Would the answer be B?